OpenStudy (anonymous):

What are a,b and c if\[\lim_{x \rightarrow 0}(\frac{ ae ^{-x} + bcos(x) + ce ^{x} }{ xsin(x) })\] is equal to 2.

OpenStudy (anonymous):

a must be 0 - first of all

OpenStudy (anonymous):

Otherwise one will have + or - infty/0 and it diverges

OpenStudy (anonymous):

a=0

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

You want any hint?

OpenStudy (anonymous):

Ok ok got it now

OpenStudy (anonymous):

take 2-nd degree taylor expansion of all three terms above the frac. line and require that the 0-th order and 1-st order combinations vanish (in the denumerat. of course)

OpenStudy (anonymous):

0-th order has the coeffs a + c +b = 0

OpenStudy (anonymous):

ok keep it up.

OpenStudy (anonymous):

1-st order requirement is -a + c = 0

OpenStudy (anonymous):

soo 2a + b = 0 ===> b = -0.5 a

OpenStudy (anonymous):

Vai ze fird oder torms du not mattah , u ahsk ?

OpenStudy (anonymous):

In english

OpenStudy (anonymous):

Vell zey duh noht mahtah since the DENOMINATOR TAylor series is of 2-nd order

OpenStudy (anonymous):

what is the answer finally?

OpenStudy (anonymous):

More precisely - the Biggest terms which will define the size of denominator are x*x because sin x ~ x near 0

OpenStudy (anonymous):

Any combination such that a = c uhnd b = -2a

OpenStudy (anonymous):

Exactly, You are smart. THANKS

OpenStudy (anonymous):

no bettah anser vill be a Medal

OpenStudy (anonymous):

actually a = 1 uhnd b = -2 iz VEE AHNSAH !

OpenStudy (anonymous):

Vich lanvige duh uou mien , eggzaktlee ?

OpenStudy (anonymous):

thx 4 z mdl @Zekarias

OpenStudy (anonymous):

tx @experimentX